# TTT: A Temporal Refinement Heuristic for Tenuously Tractable Discrete Time Reachability Problems

### v1

@@ 6.1 Results
In fig. 3, one can observe that the approximate error generally decreases with increasing budget. Note that the trend is not monotonic because the algorithm is a heuristic.
We compare our algorithm to the hand tuned hybrid symbolic approach used in [8] which is represented in  fig. 3 as dotted lines. One might expect that the hand tuned approach would produce approximate reachable sets more quickly for a given amount of error. A significant amount of work was presumably done offline to identify good symbolic intervals that must instead be identified at runtime with our approach. However, it turns out that our heuristic can actually produce approximate reachable sets with similar amounts of error as the hand-tuned approach in 20-70% less time, as shown in fig. 3. Furthermore, our approach is able to produce reachable sets for some problems such as the pendulum problem S1 that have 40% less error than the hand-tuned approach when allowed to compute for longer than the time taken by the hand-tuned approach.
Figure 3: Approximate reachable sets of various fidelity are produced depending on allotted time and are competitive with hand-tuned baseline indicated with dashed lines.
We also compare sets produced by our algorithm to concrete one step reachable sets. Using one step sets is fast but leads to large approximation error. Examples of the reachable sets computed by our algorithm and baselines can be seen in  fig. 4 and fig. 5. Set are compared both using eq. 4 as well as the ratio of the sum of radii:
[…]
One step concrete sets are labeled ’naive’ and one can observe that every set produced by our approach is equal to or tighter than those produced with the naive approach. To compare to the hand-tuned approach, we select a run from fig. 3 that had roughly equivalent or less error in a smaller amount of time and calculate the increase in speed by our approach. For the runs selected in fig. 4 and fig. 5, not every set produced by our approach is tighter than sets produced by the hand tuned approach, but when error is calculated over all sets in the trajectory, the total error is roughly equal or smaller. For pendulum problem S1 featured fig. 4, our algorithm is 20.9% faster than the hand-tuned approach for less error by both radius and volume metrics, and for the TORA problem T1 featured in fig. 5, our algorithm is 24% faster for an $\approx$ equal amount of error (1.5% less error according to the ratio metric and 1.2% more error according to the volume metric).
In fig. 6, we display numerical results for the remaining problems, car problems C1 and C2. For C1 and C2, we select runs that have less error by both radius and volume metrics and are respectively 70.8% faster and 65.2% faster, which translates to 4.7 minutes faster and 4.1 minutes faster.
Naive Hand-Tuned Ours
Error (volume) 11.9 2.27 1.94
Error (radius) 3.79 1.61 1.59
Time 7.84s 19.3s 15.3s
Figure 4: Reachable sets for pendulum problem S1 with two layer, 25 neuron-per-layer neural network controller show less error than the naive baseline and greater speed than the hand-tuned baseline for less error.
Naive Hand-Tuned Ours
Error (volume) 31.7 7.67 7.78
Error (radius) 1.81 1.41 1.38
Time 6.62s 63.0s 47.8s
Figure 5: Dimensions 1 and 2 of the reachable sets for TORA problem T1 with three layer, 25 neuron-per-layer neural network controller which has less error than the naive baseline and greater speed than the hand-tuned baseline for similar error.
Car (C1) One-Step Hand-Tuned Ours
Error (volume) 37.7 4.90 3.89
Error (radius) 1.88 1.31 1.28
Time 10.3s 400s 117s
Car (C2)
Error (volume) 37.9 5.29 4.43
Error (radius) 1.85 1.30 1.29
Time 11.1s 378s 131s
Figure 6: Reachable sets for the Car problems C1 and C2 are computed significantly faster for comparable error using our algorithm as compared to the hand-tuned baseline.
Lastly, our code will be released as an open source tool and be available here.
6.2 Limitations
The approach described here has limitations related to budgeting. A poor choice of initial solution from optimizer can lead to budget overruns as the algorithm will continue to compute if a finite solution has not yet been found. Further, budgeting with the linear time approximation as described is not accurate for all reachability problems, leading to budget overruns and/or conservative solutions for some problems; which may be exacerbated by long horizon queries. As previously mentioned, our algorithm is a heuristic. There does not exist a guarantee that given two budgets $b_{A}>b_{B}$ the sets produced with budget $b_{A}$ will be tighter than those produced using $b_{B}$.
7 Conclusion
In this paper we have introduced an algorithm for temporal refinement of tenuously tractable discrete time reachability problems. The algorithm computes reachable sets using constraints from several timesteps at once in order to produce tighter sets. The temporal depth of timestep constraints considered for a given reachable set is calculated based upon remaining computational budget and estimated future computation. The algorithm introduced here makes it possible to do hybrid-symbolic reachability without hand tuning concrete and symbolic intervals. We demonstrated this algorithm on a difficult class of reachability problems: nonlinear dynamical systems with neural network control policies, and demonstrated that it is able to produce reachable sets of varying error given varying computational budget. Additionally, it is able to produce reachable sets 20-70% faster then hand tuned approaches for the same amount of error. Ultimately, this algorithm and associated open-source code represent tools to enable more efficient reachability analysis for complex control systems.

### v2

@@ V Soundness, Completeness, Complexity
Lemma V.2
Alg.1 will terminate producing $n$ reachable sets ~[~$\tilde{\mathbb{R}}\triangleq[\tilde{\mathcal{R}}_{1},\ldots,\tilde{\mathcal{R}}_{n}]$ under the assumption that the true reachable set is non-emtpy.
Proof:
All input variables are bounded and therefore each MIP is guaranteed to terminate in finite time when early stopping is not used. When early stopping is used, the algorithm will extend the compute time until a finite feasible solution with $<50\%$ relative duality gap has been found. Such a solution is guaranteed to exist and to be found in finite time. $\square$
Theorem V.3
Alg. 1 performs hybrid-symbolic reachability analysis with an automatically chosen mixture of concrete and symbolic queries to produce $n$ overapproximate reachable sets ~[~$\tilde{\mathbb{R}}\triangleq[\tilde{\mathcal{R}}_{1},\ldots,\tilde{\mathcal{R}}_{n}]$.
Proof:
See lemmas 5.1 and 5.2. $\square$
We note that in addition to producing $n$ overapproximate reachable sets, Alg. 1 also heuristically minimizes overapproximation error and heuristically adheres to the budget.
Next, we reason about the behavior of Alg. 1 in the limit of infinite budget in Corollary V.5 and the worst-case complexity of Alg. 1 in Corollary LABEL:col:complex.
Definition V.4
The tightest sequence of approximate reachable sets $\tilde{\mathbb{R}}|$ possible given a fixed, separable [chen2023one] subroutine symbolic_reach is the sequence computed with pure symbolic reachability wherein each query begins from $t=0,~{}\mathcal{X}_{0}$ and extends to the current time $t$ for $t\in\{1,\ldots,n\}$.
Corollary V.5
In the limit when the time budget $b\rightarrow\infty$, the algorithm will produce $\tilde{\mathbb{R}}|$.